OpenStudy (anonymous):

3x-2y=12 and 5x+2y=4 solve using the addition method

OpenStudy (anonymous):

2 and -3

OpenStudy (anonymous):

(3x-2y)+(5x+2y)=12+4 8x=16 x=2 y=-3

OpenStudy (anonymous):

thanks avadh